# Cross-Modal Reconstruction Pretraining for Ramp Flow Prediction at Highway Interchanges

### v1

@@ 3.3 Downstream prediction model
Specifically, a long sequence of $T_{\text{long}}$ time steps is first fed into the pretrained spatial and temporal encoders to generate the corresponding spatial representation $H^{(S)}$ and temporal representation $H^{(T)}$. Subsequently, the commonly used short-term input sequence $X_{t-(T-1):t}$ is fed into the downstream prediction model to obtain the hidden representation $H^{(F)}\in\mathbb{R}^{T\times M\times D}$, where $D$ is the dimensionality of the hidden representation. During the input stage, the same masking strategy as in the pretraining stage is applied to ensure consistency.
To align with the hidden representation $H^{(F)}$, we extract the last $T^{\prime}$ patches from $H^{(S)}$ and $H^{(T)}$ and reshape them into tensors $H^{\prime(S)}\in\mathbb{R}^{T\times M\times D}$ and $H^{\prime(T)}\in\mathbb{R}^{T\times M\times D}$, respectively. These representations are then projected to the target dimension $D^{\prime}$ using two independent two-layer multilayer perceptrons.
The final augmented representation $H^{(\text{Aug})}\in\mathbb{R}^{T\times M\times D}$ is obtained by summing the projected spatial and temporal representations with the downstream hidden representation, as follows:
$H^{(\text{Aug})}=\text{MLP}(H^{\prime(S)})+\text{MLP}(H^{\prime(T)})+H^{(F)}.$ (7)
Up to this point, the augmented representation $H^{(\text{Aug})}$ incorporates the short-term representation learned by the downstream prediction model itself and the long temporal dependency features provided by STDAE, thus effectively compensating for the limited ability of traditional models to capture long-term dependencies and significantly improving the overall performance and generalization of downstream spatio-temporal prediction tasks.
In this study, the downstream forecasting model adopts GWNet (Wu et al., 2019). STDAE learns structured global representations by separately modeling spatial and temporal features, without requiring additional masking or reconstruction steps. By integrating the STDAE representations into the hidden states as auxiliary information, the model can better capture long-range dependencies, enabling more accurate and stable predictions in complex spatiotemporal tasks.
@@ 4 Dataset and processing
This section provides a detailed description of the entire process of data collection, cleaning, and feature engineering used in this study. The overall workflow is illustrated in Figure 2. The purpose of this process is to transform raw ETC transaction records into high-quality time series data of mainline and ramp traffic states that can be utilized by the model.
@@ 4.3 Data cleaning
The raw ETC transaction records are discrete event logs, which cannot be directly used for time series analysis. Therefore, the following data processing and feature engineering procedures are designed to transform the raw records into structured traffic features. First, the format of the acquired data is validated to ensure that all records follow a consistent structure and contain complete fields, which lays the foundation for automated processing. Erroneous records, such as timestamps outside the valid range or duplicate records of the same vehicle within the same second, are then removed. Next, each ETC record is matched with the corresponding gantry information using the gantry ID as the key index so that accurate geographical and road segment attributes can be assigned. Afterward, continuous event data are aggregated into discrete time series by defining the sampling granularity and splitting the study period into time windows. Mainline traffic volumes are computed by counting the number of vehicles passing through each gantry during every time window. To estimate mainline speed, the same license plate is matched across two adjacent gantries, the vehicle travel time over the segment is calculated, and the distance between gantries (obtained from kilometer markers) is divided by the travel time. The arithmetic mean of all vehicle speeds within a window is taken as the segment’s average speed. Finally, ramp flows, which are a key feature in this study, are inferred by matching license plates observed at upstream and downstream gantries, since ramps typically lack ETC gantries. Through this complete process, three core time series feature sets are obtained: mainline volume, mainline speed, and ramp volume.
In this study, the data covers the period from September 7, 2024, to September 29, 2024, totaling 23 days. Sampling intervals of 3 minutes, 5 minutes, and 10 minutes are adopted. A z-score normalization method is applied. The dataset is split into training, validation, and test sets in a ratio of 17:3:3, where the first 17 days are used as the training set, the middle 3 days as the validation set, and the last 3 days as the test set.
@@ 5 Experiments and results
This section presents a comprehensive evaluation of the proposed STDAE model on the datasets from Section 4. GWNet is selected as the downstream prediction model (Wu et al., 2019), forming the combined STAEGWNET model. Firstly, the overall performance of the STAEGWNET model is evaluated through comparison with several state-of-the-art models. Subsequently, ablation experiments are conducted to validate the effectiveness of the pretraining approach, followed by a deep analysis of the experimental results.
@@ 5.1 Experimental setup
Baselines. We compare STAEGWNET with four groups of representative benchmark models. The time-series models include GRU and TCN. The spatio-temporal GNN models comprise AGCRN (Bai et al., 2020), DCRNN (Li et al., 2017), MTGNN (Wu et al., 2020b), STPGNN (Kong et al., 2024), TGCN (Zhao et al., 2019), STGCN (Yu et al., 2018), and D2STGNN (Shao et al., 2022b). The spatio-temporal enhanced CNN models include GWNet (Wu et al., 2019) and STNorm (Deng et al., 2021). The spatio-temporal Transformer models consist of STAEformer (Liu et al., 2023a) and iTransformer (Liu et al., 2024a). These baselines span traditional statistical and machine learning methods to the latest graph neural networks and Transformer-based architectures, providing a comprehensive evaluation of the spatio-temporal modeling and forecasting capabilities of STAEGWNET. The input to these baseline models is consistent with STAEGWNET, consisting of mainline feature time series. In addition, we compare the results of HimNet (Dong et al., 2024), GWNet (Wu et al., 2019), and STDMAE (Gao et al., 2024) when using ramp flow time series as input with our model using mainline feature time series as input. The purpose is to verify whether using mainline features as input can achieve better prediction performance than directly using ramp flow sequences, thereby demonstrating the feasibility of predicting using only mainline data.
Settings. In the pretraining stage, the long temporal horizon is set to $T_{\text{long}}=1day$. In the forecasting stage, the input sequence length is fixed at $T=12$ time steps (Gao et al., 2024; Song et al., 2020) and the embedding dimension is $G=96$. The encoder comprises four Transformer layers and the decoder one Transformer layer, both employing four-head multi-head attention. To match the forecasting input, the patch size is $L=12$, and we set $T^{\prime}=1$, retaining only the last patch from the spatial representation $H^{(S)}$ and temporal representation $H^{(T)}$. Training uses the Adam optimizer with an initial learning rate of 0.002 and the mean absolute error (MAE) loss. Performance is evaluated by mean absolute error (MAE), mean absolute percentage error (MAPE), and root mean squared error (RMSE). The results are averaged over 10 independent runs. All experiments are conducted on a Windows workstation equipped with a single NVIDIA V100 Tensor Core GPU. The experiments are conducted using the platform (Shao et al., 2024).
@@ 5.2 Overall performance
The performance of models is listed in Table 5, Table 5, and Table 5. The proposed STDAEGWNET model demonstrates significant superiority across all three interchange datasets. Moreover, for the 3-minute, 5-minute, and 10-minute forecasting tasks, the model achieves the best or second-best performance on the MAE, MAPE, and RMSE metrics, consistently maintaining the top overall average ranking.
Specifically, the model performance is reflected in the following four aspects. Overall Best Performance. The average ranking results further validate the comprehensive superiority of STDAEGWNET. The model average ranks on the QiLin, DanYangXinQu, and XueBu datasets are 1.00, 2.11, and 1.44, respectively, all ranking first. This performance clearly surpasses existing mainstream models, including graph neural network models, whose best ranks on the three datasets reach only 2.11, and Transformer-based models, whose best rank is only 4.22. This demonstrates that the proposed pretraining and downstream prediction framework can more effectively capture the spatiotemporal dependencies between mainline and ramp traffic, thereby exhibiting outstanding performance in practical prediction tasks. Significant Accuracy Advantage. In most prediction scenarios, STDAEGWNET achieves the lowest MAE, indicating that its predictions deviate least from the true values and outperform other baseline models in accuracy. At a 3 min sampling interval, the model mean absolute errors on the QiLin, DanYangXinQu, and XueBu datasets are 4.89, 5.61, and 4.58, respectively, all the lowest among the compared models and statistically significant. Even in scenarios where it does not achieve absolute optimality, performance remains close to the best, maintaining stable competitiveness. Strong Stability and Robustness. The root mean square error results show that STDAEGWNET generally outperforms other models, indicating concentrated prediction distributions and fewer extreme errors. On the QiLin dataset for 5 min predictions, STDAEGWNET achieves a root mean square error of 11.29, significantly lower than GRU at 13.90, TCN at 13.91, and DCRNN at 12.36. On the XueBu dataset for 10 min predictions, root mean square error is 23.70, lower than D2STGNN at 24.15 and GWNet at 24.18, demonstrating the model’s stability and robustness. Good Generalizability. Across three different hub datasets, STDAEGWNET consistently ranks among the top, indicating strong cross-scenario generalization and adaptability to diverse traffic characteristics. Compared with models that excel only on a single dataset, STDAEGWNET shows more robust overall performance. D2STGNN performs strongly on the XueBu and DanYangXinQu datasets with average ranks of 2.11 and 2.56 but performs relatively weaker on the QiLin dataset at 3.56. In contrast, STDAEGWNET maintains the best average ranks across all three datasets at 1.00, 2.11, and 1.44, highlighting superior robustness and consistency across scenarios.
@@
This comparative experiment highlights a central and important insight. Mainline traffic features contain richer and more predictive contextual information than the ramp’s own historical sequence. This seemingly counterintuitive result that indirect data can outperform direct data can be explained by the fact that ramp traffic flow largely reflects the outcome or manifestation of the broader traffic dynamics within the network. In contrast, mainline features such as flow and speed more effectively capture the underlying causes of ramp flow variations, such as upstream arrival waves or downstream capacity constraints. The STDAEGWNET model successfully learns these fundamental traffic evolution patterns from mainline data, which are indirect yet more informative, enabling it to deliver more accurate predictions than local perspective models that rely solely on the ramp’s own history.
In summary, the experimental findings strongly validate the effectiveness of the STDAEGWNET framework. It provides a practical solution to the challenge of unavailable ramp historical data under real-time blind spot conditions, and it achieves prediction accuracy that rivals or even surpasses advanced models relying on direct ramp history, thereby demonstrating dual advantages in both theory and practice.
@@ 5.3 Ablation study
Ablation Study on the Encoding Mechanism. To assess the effectiveness of the spatiotemporal encoding strategy, we design three model variants:
@@
Figure 3: Comparison of STDAE and its ablated versions on three datasets with 5-minute sampling granularity.
We report the experimental results on three datasets with a sampling granularity of 5 min. As shown in Figure 3, STDAE consistently outperforms all ablated versions. This advantage mainly stems from its design of reconstructing ramp traffic flow by leveraging mainline traffic features. Compared with approaches that directly rely on historical ramp flow, this mechanism extracts richer contextual information from mainline data, thereby compensating for the absence of ramp information under “real-time blind spot” conditions. In the spatial dimension, mainline flow reflects the traffic environment upstream and downstream of the ramp. In the temporal dimension, mainline features capture the underlying causes of ramp flow fluctuations, such as upstream arrival waves and downstream capacity constraints. Through this cross-source reconstruction mechanism, STDAE is able to learn the evolutionary patterns between ramps and mainlines, thereby significantly enhancing prediction accuracy and robustness.
Ablation Study on Predictor. To evaluate the generality of STDAE, we test five downstream predictors with different backbones, including RNN, GCN+RNN, GCN+TCN, and Transformer:
@@
    STDAE-GWNET: Using GWNet as the predictor.
The experiments were conducted on three datasets with a 5 min sampling granularity. Table 7 presents the ablation results of STDAE across different downstream predictors. The experiments show that for all five downstream predictors, models enhanced with STDAE exhibit consistent and significant performance improvements across the three datasets.
@@
STDAE-GWNET $\mathbf{7.01}^{*}/\mathbf{11.29}^{*}/\mathbf{0.54}$ $\mathbf{8.59}^{*}/\mathbf{19.57}/\mathbf{0.53}$ $\mathbf{6.89}^{*}/\mathbf{12.89}^{*}/\mathbf{0.52}^{*}$
@@ 5.4 Robustness on Missing Data
To evaluate the performance of the STDAEGWNET model under missing mainline data conditions, we introduced directional and temporal masking strategies to the mainline data to simulate data loss. The directional mask hides data from two mainline ETC gantries in the eastbound direction, while the temporal mask hides the last 6 steps out of every 12 time steps. Experiments were conducted on three datasets with sampling intervals of 3 min, 5 min, and 10 min. We compared models with and without STDAE, and the MAE results are shown in Figure 4. The results indicate that, across all datasets and sampling intervals, models incorporating STDAE significantly improve prediction accuracy. Compared to baseline models without STDAE, the STDAE-enhanced models consistently demonstrate clear advantages, achieving an average MAE reduction of approximately 2.23%. Moreover, the majority of the results passed the significance test ($p<0.01$).
@@
Figure 5: Ramp-level Prediction of the QiLin Interchange with a 5 min Sampling Interval. (a) step = 3. (b) step = 6. (c) step = 12.

### v2

@@ 3.3 Downstream prediction model
Specifically, a long sequence of $T_{\text{long}}$ time steps is first fed into the pretrained spatial and temporal encoders to generate the corresponding spatial representation $H^{(S)}$ and temporal representation $H^{(T)}$. Subsequently, the commonly used short-term input sequence $X_{t-(T-1):t}$ is fed into the downstream prediction model to obtain the hidden representation $H^{(F)}\in\mathbb{R}^{T\times M\times D}$, where $D$ is the dimensionality of the hidden representation. During the input stage, the same masking strategy as in the pre-training stage is applied to ensure consistency.
To align with the hidden representation $\mathbf{H}^{(F)}$, the last $T^{\prime}$ patches from $\mathbf{H}^{(S)}$ and $\mathbf{H}^{(T)}$ are extracted and reshaped into tensors $\mathbf{H}^{\prime(S)}\in\mathbb{R}^{T\times M\times D}$ and $\mathbf{H}^{\prime(T)}\in\mathbb{R}^{T\times M\times D}$, respectively. These representations are then projected to the target dimension $D^{\prime}$ using two independent two-layer multilayer perceptrons.
The final augmented representation $\mathbf{H}^{(\text{Aug})}\in\mathbb{R}^{T\times M\times D}$ is obtained by summing the projected spatial and temporal representations with the downstream hidden representation, as follows:
$\mathbf{H}^{(\text{Aug})}=\text{MLP}(\mathbf{H}^{\prime(S)})+\text{MLP}(\mathbf{H}^{\prime(T)})+\mathbf{H}^{(F)}.$ (22)
Up to this point, the augmented representation $\mathbf{H}^{(\text{Aug})}$ incorporates the short-term representation learned by the downstream prediction model itself and the long temporal dependency features provided by STDAE, thus effectively compensating for the limited ability of traditional models to capture long-term dependencies and significantly improving the overall performance and generalization of downstream spatio-temporal prediction tasks.
In this study, the downstream forecasting model adopts GWNet [ijcai2019p264]. STDAE learns structured global representations by separately modeling spatial and temporal features, without requiring additional masking or reconstruction steps. By integrating the STDAE representations into the hidden states as auxiliary information, the model can better capture long-range dependencies, enabling more accurate and stable predictions in complex spatiotemporal tasks.
@@ 4 Dataset and processing
This section provides a detailed description of the entire process of data collection, cleaning, and feature engineering used in this study. The overall workflow is illustrated in Figure 3. The purpose of this process is to transform raw ETC transaction records into time series data of mainline and ramp traffic states that can be utilized by the model.
@@ 4.3 Data cleaning
The raw ETC transaction records are discrete event logs, which cannot be directly used for time series analysis. Therefore, the following data processing and feature engineering procedures are designed to transform the raw records into structured traffic features. First, the format of the acquired data is validated to ensure that all records follow a consistent structure and contain complete fields, which lays the foundation for automated processing. Erroneous records, such as timestamps outside the valid range or duplicate records of the same vehicle within the same second, are then removed. Next, each ETC record is matched with the corresponding gantry information using the gantry ID as the key index so that accurate geographical and road segment attributes can be assigned. Afterward, continuous event data are aggregated into discrete time series by defining the sampling intervals and splitting the study period into time windows. Mainline traffic volumes are computed by counting the number of vehicles passing through each gantry during every time window. To estimate mainline speed, the same license plate is matched across two adjacent gantries, the vehicle travel time over the segment is calculated, and the distance between gantries (obtained from kilometer markers) is divided by the travel time. The arithmetic mean of all vehicle speeds within a window is taken as the segment’s average speed. Finally, ramp flows, which are a key feature in this study, are inferred by matching license plates observed at upstream and downstream gantries, since ramps typically lack ETC gantries. Through this complete process, three core time series feature sets are obtained: mainline volume, mainline speed, and ramp volume.
In this study, the data covers the period from September 7, 2024, to September 29, 2024, totaling 23 days. After data cleaning and preprocessing, a total of 3,962,597 valid vehicle passage records are obtained. Sampling intervals of 3 min, 5 min [WOS:001347142804014, shao2024exploring], and 10 min are adopted. Multiple sampling intervals are employed to assess the model’s performance under varying temporal granularities. A z-score normalization method is applied. The dataset is split into training, validation, and test sets in a ratio of 17:3:3, where the first 17 days are used as the training set, the middle 3 days as the validation set, and the last 3 days as the test set.
@@ 5 Experiments and results
This section presents a comprehensive evaluation of the proposed STDAE model on the datasets from Section 4. GWNet is selected as the downstream prediction model [ijcai2019p264], forming the combined STDAEGWNET model. Firstly, the overall performance of the STDAEGWNET model is evaluated through comparison with several state-of-the-art models. Subsequently, ablation experiments are conducted to validate the effectiveness of the pre-training approach, followed by a deep analysis of the experimental results.
@@ 5.1 Experimental setup
Baselines. In this section, STDAEGWNET is compared with four groups of representative benchmark models. The time-series models include GRU and TCN. The spatio-temporal GNN models comprise AGCRN [bai2020adaptive], DCRNN [li2017diffusion], MTGNN [wu2020connecting], STPGNN [kong2024spatio], TGCN [zhao2019t], STGCN [WOS:000764175403107], and D2STGNN [shao2022decoupled]. The spatio-temporal enhanced CNN models include GWNet [ijcai2019p264] and STNorm [deng2021st]. The spatio-temporal Transformer models consist of STAEformer [liu2023spatio] and iTransformer [liu2024itransformer]. These baselines span traditional statistical and machine learning methods to the latest graph neural networks and Transformer-based architectures, providing a comprehensive evaluation of the spatio-temporal modeling and forecasting capabilities of STDAEGWNET. The input to these baseline models is consistent with STDAEGWNET, consisting of mainline feature time series. In addition, the performance of HimNet [dong2024heterogeneity], GWNet [ijcai2019p264], and STDMAE [WOS:001347142804014] with ramp flow time series as input is compared to that of our model using mainline feature time series. The purpose is to verify whether using mainline features as input can achieve better prediction performance than directly using ramp flow sequences, thereby demonstrating the feasibility of predicting using only mainline data.
Settings. In the pre-training stage, the long temporal horizon is set to $T_{\text{long}}=1day$. In the forecasting stage, the input sequence length is fixed at $T=12$ time steps [WOS:001347142804014, song2020spatial] and the embedding dimension is $G=96$. The encoder comprises four Transformer layers and the decoder one Transformer layer, both employing four-head multi-head attention. To match the forecasting input, the patch size is $L=12$, and $T^{\prime}=1$, retaining only the last patch from the spatial representation $\mathbf{H}^{(S)}$ and temporal representation $\mathbf{H}^{(T)}$. Training uses the Adam optimizer with an initial learning rate of 0.002 and the mean absolute error (MAE) loss. Performance is evaluated by mean absolute error (MAE), mean absolute percentage error (MAPE), and root mean squared error (RMSE). The results are averaged over 10 independent runs. All experiments are conducted on a Windows workstation equipped with a single NVIDIA V100 Tensor Core GPU. The experiments are conducted using the platform [shao2024exploring].
@@ 5.2 Overall performance
The performance of models is listed in Table 5, Table 5, and Table 5. The proposed STDAEGWNET model demonstrates significant superiority across all three interchange datasets. Moreover, for the 3 min, 5 min, and 10 min forecasting tasks, the model achieves the best or second-best performance on the MAE, MAPE, and RMSE metrics, consistently maintaining the top overall average ranking.
Specifically, the model performance is reflected in the following four aspects. Overall Performance. The average ranking results further validate the comprehensive superiority of STDAEGWNET. The model average ranks on the QiLin, DanYangXinQu, and XueBu datasets are 1.00, 2.11, and 1.44, respectively, all ranking first. This performance clearly surpasses existing mainstream models, including graph neural network models, whose best ranks on the three datasets reach only 2.11, and Transformer-based models, whose best rank is only 4.22. This demonstrates that the proposed pre-training and downstream prediction framework can more effectively capture the spatiotemporal dependencies between mainline and ramp traffic, thereby exhibiting outstanding performance in practical prediction tasks. Accuracy. In most prediction scenarios, STDAEGWNET achieves the lowest MAE, indicating that its predictions deviate least from the true values and outperform other baseline models in accuracy. At a 3 min sampling interval, the model mean absolute errors on the QiLin, DanYangXinQu, and XueBu datasets are 4.89, 5.61, and 4.58, respectively, all the lowest among the compared models and statistically significant. Even in scenarios where it does not achieve absolute optimality, performance remains close to the best, maintaining stable competitiveness. Stability and Robustness. The RMSE results show that STDAEGWNET generally outperforms other models, indicating concentrated prediction distributions and fewer extreme errors. On the QiLin dataset for 5 min predictions, STDAEGWNET achieves a RMSE of 11.29, significantly lower than GRU at 13.90, TCN at 13.91, and DCRNN at 12.36. On the XueBu dataset for 10 min predictions, root mean square error is 23.70, lower than D2STGNN at 24.15 and GWNet at 24.18, demonstrating the model’s stability and robustness. Generalizability. Across three different hub datasets, STDAEGWNET consistently ranks among the top, indicating strong cross-scenario generalization and adaptability to diverse traffic characteristics. Compared with models that excel only on a single dataset, STDAEGWNET shows more robust overall performance. D2STGNN performs strongly on the XueBu and DanYangXinQu datasets with average ranks of 2.11 and 2.56 but performs relatively weaker on the QiLin dataset at 3.56. In contrast, STDAEGWNET maintains the best average ranks across all three datasets at 1.00, 2.11, and 1.44, highlighting superior robustness and consistency across scenarios.
@@
Figure 4: Comparison of MAPE across different sampling intervals. (a) Average MAPE of 14 models. (b) MAPE of the STDAEGWNET model.
The impact of sampling interval on model prediction performance is investigated by comparing the models on three interchanges—QiLin, DanYangXinQu, and XueBu—using 3 min, 5 min, and 10 min sampling intervals, with performance evaluated using MAPE, as shown in Figure 4. The average performance of the models in Tables 5, 5, and 5 shows that for the QiLin dataset, MAPE decreases as the sampling interval increases, from 65.5% at 3 min to 57.2% at 10 min, indicating that the 10 min interval is optimal. In contrast, for the DanYangXinQu dataset, MAPE increases from 57.7% at 3 min to 64.2% at 10 min, making the 3 min interval the best. The XueBu dataset performs best at a 5 min interval, with MAPE of 58.5%, lower than 59.2% at 3 min and 69.8% at 10 min. Taking the STDMAEGWNET model as an example, its optimal MAPE values on the QiLin, DanYangXinQu, and XueBu datasets are 46.8% at 10 min, 51.4% at 3 min, and 51.6% at 5 min, respectively, further confirming this trend. This variation likely results from differences in traffic flow fluctuations and trends across the interchanges: QiLin exhibits a clear trend, making longer intervals beneficial for smoothing noise; DanYangXinQu experiences frequent traffic fluctuations, where shorter intervals help retain critical dynamic information; XueBu represents an intermediate case, with a 5 min interval balancing noise smoothing and information retention. Therefore, selecting an appropriate sampling interval for each interchange is an important factor in improving model prediction performance.
@@ 5.3 Ablation study
The effectiveness and necessity of the proposed spatiotemporal encoding strategy are evaluated through three carefully designed model variants, which separately examine the roles of temporal and spatial dependencies. These variants help clarify how each component contributes to the overall reconstruction capability and predictive performance of STDAE.
@@
Figure 5: Comparison of STDAE and its ablated versions on three datasets with 5 min sampling interval.
The experimental results for the three datasets, using a 5 min sampling interval, are presented in Figure 5. STDAE consistently outperforms all ablated versions. This advantage mainly stems from its design of reconstructing ramp traffic flow by leveraging mainline traffic features. Compared with approaches that directly rely on historical ramp flow, this mechanism extracts richer contextual information from mainline data, thereby compensating for the absence of ramp information under “real-time blind spot” conditions. In the spatial dimension, mainline flow reflects the traffic environment upstream and downstream of the ramp. In the temporal dimension, mainline features capture the underlying causes of ramp flow fluctuations, such as upstream arrival waves and downstream capacity constraints. Through this cross-source reconstruction mechanism, STDAE is able to learn the evolutionary patterns between ramps and mainlines, thereby significantly enhancing prediction accuracy and robustness.
The generality of the proposed STDAE framework are further examined by integrating it with five downstream predictors featuring diverse architectural backbones, including RNN, GCN+RNN, GCN+TCN, and Transformer:
@@
    STDAEGWNET: Using GWNet as the predictor.
The experiments are conducted on three datasets with a 5 min sampling interval. Table 7 presents the ablation results of STDAE across different downstream predictors. The experiments show that for all five downstream predictors, models enhanced with STDAE exhibit consistent and significant performance improvements across the three datasets.
@@
STDAEGWNET $\mathbf{7.01}^{*}/\mathbf{11.29}^{*}/\mathbf{0.54}$ $\mathbf{8.59}^{*}/\mathbf{19.57}/\mathbf{0.53}$ $\mathbf{6.89}^{*}/\mathbf{12.89}^{*}/\mathbf{0.52}^{*}$
@@ 5.4 Robustness on Missing Data
To evaluate the performance of the STDAEGWNET model under missing mainline data conditions, directional and temporal masking strategies are applied to the mainline data to simulate data loss. The directional mask hides data from two mainline ETC gantries in the eastbound direction, while the temporal mask hides the last 6 steps out of every 12 time steps. Experiments are conducted on three datasets with sampling intervals of 3 min, 5 min, and 10 min. Models with and without STDAE are compared, and the MAE results are shown in Figure 6. The results indicate that, across all datasets and sampling intervals, models incorporating STDAE significantly improve prediction accuracy. Compared to baseline models without STDAE, the STDAE-enhanced models consistently demonstrate clear advantages, achieving an average MAE reduction of approximately 2.23%. Moreover, the majority of the results passed the significance test ($p<0.01$).
@@
Figure 7: Ramp-level Prediction of the QiLin Interchange with a 5 min Sampling Interval (step = 3). Subfigure labels correspond to the source direction: (a) East, (b) West, (c) South, and (d) North. Turning movements are indicated as follows: (1) right-turn, (2) straight, and (3) left-turn. Various traffic dynamics: (a4) an upward trend, (b4) a fluctuating trend, (c4) a downward trend, and (d4) a fluctuating upward trend.